15 km² - 52 ha =
47 m² - 3207 cm² =
1086 cm² / 19 mm² =

Answer by checkley71(8403)

15km^2-52ha no factoring.
47m^2-3207cm^2=m^2(47-3207c)
1086cm^2/19mm^2=1086c/19m